# On Selecting Few-Shot Examples for LLM-based Code Vulnerability Detection

### v1

@@ Appendix C Different Variants of Learn-from-Mistakes (LFM)
Table 3. Performance comparison of different LFM variations. Each cell shows the mean and standard deviation over five runs. For the LFM variation names, the following abbreviations are used: S = stacked, U = unstacked, 1× = one iteration, m× = more than one iteration, inc. = incorrect, corr. = correct, and gray = gray cases. ‘–’ indicates metrics that are undefined due to division by zero. Dataset LFM Variations Gemma-3-4b-it Qwen-2.5-Coder Accuracy Precision Recall F1 Score Accuracy Precision Recall F1 Score DiverseVul S-1× (inc.) 0.525 (±0.003) 0.840 (±0.056) 0.061 (±0.005) 0.114 (±0.009) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.502 (±0.013) 0.501 (±0.007) 0.940 (±0.025) 0.654 (±0.010) 0.502 (±0.016) 0.501 (±0.010) 0.844 (±0.063) 0.628 (±0.022) U-m× (inc.) 0.507 (±0.010) 0.504 (±0.005) 0.995 (±0.008) 0.669 (±0.003) 0.527 (±0.039) 0.516 (±0.025) 0.976 (±0.045) 0.674 (±0.009) U-m× (corr.) 0.597 (±0.053) 0.778 (±0.056) 0.295 (±0.189) 0.383 (±0.204) 0.621 (±0.065) 0.752 (±0.027) 0.365 (±0.199) 0.455 (±0.199) U-m× (gray) 0.513 (±0.000) 0.667 (±0.000) 0.053 (±0.000) 0.099 (±0.000) 0.598 (±0.061) 0.713 (±0.059) 0.341 (±0.226) 0.415 (±0.199) NodeMedic S-1× (inc.) 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.720 (±0.000) 0.723 (±0.000) 0.993 (±0.000) 0.837 (±0.000) U-1× (inc.) 0.273 (±0.004) 0.500 (±0.707) 0.001 (±0.003) 0.003 (±0.006) 0.274 (±0.005) 0.333 (±0.422) 0.004 (±0.006) 0.009 (±0.011) U-m× (inc.) 0.277 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.272 (±0.003) 0.000 (±0.000) 0.000 (±0.000) U-m× (corr.) 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.724 (±0.002) 0.727 (±0.001) 0.991 (±0.003) 0.839 (±0.001) U-m× (gray) 0.506 (±0.003) 0.765 (±0.005) 0.460 (±0.000) 0.574 (±0.001) 0.467 (±0.100) 0.748 (±0.022) 0.412 (±0.245) 0.493 (±0.158) PrimeVul S-1× (inc.) 0.516 (±0.013) 0.726 (±0.186) 0.437 (±0.459) 0.336 (±0.270) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.507 (±0.010) 0.529 (±0.054) 0.773 (±0.352) 0.551 (±0.212) 0.497 (±0.027) 0.489 (±0.042) 0.466 (±0.137) 0.470 (±0.092) U-m× (inc.) 0.496 (±0.009) 0.397 (±0.221) 0.034 (±0.022) 0.061 (±0.038) 0.501 (±0.005) 0.502 (±0.005) 0.864 (±0.219) 0.621 (±0.077) U-m× (corr.) 0.499 (±0.002) 0.499 (±0.001) 0.886 (±0.223) 0.626 (±0.079) 0.489 (±0.010) 0.386 (±0.090) 0.034 (±0.010) 0.062 (±0.018) U-m× (gray) 0.499 (±0.002) 0.496 (±0.009) 0.366 (±0.284) 0.357 (±0.202) 0.498 (±0.017) 0.529 (±0.141) 0.246 (±0.241) 0.267 (±0.198) SVENC S-1× (inc.) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.500 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) U-m× (inc.) 0.500 (±0.000) - 0.000 (±0.000) 0.000 (±0.000) 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) U-m× (corr.) 0.504 (±0.009) 0.503 (±0.006) 0.960 (±0.069) 0.659 (±0.013) 0.493 (±0.009) 0.496 (±0.005) 0.911 (±0.056) 0.642 (±0.016) U-m× (gray) 0.478 (±0.000) 0.474 (±0.000) 0.400 (±0.000) 0.434 (±0.000) 0.502 (±0.011) 0.080 (±0.109) 0.114 (±0.142) SVENP S-1× (inc.) 0.500 (±0.000) 0.500 (±0.000) 0.905 (±0.013) 0.644 (±0.003) 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) U-1× (inc.) 0.558 (±0.067) 0.571 (±0.083) 0.600 (±0.127) 0.573 (±0.050) 0.590 (±0.051) 0.607 (±0.074) 0.547 (±0.111) 0.567 (±0.062) U-m× (inc.) 0.540 (±0.033) 0.567 (±0.079) 0.447 (±0.130) 0.483 (±0.074) 0.571 (±0.054) 0.569 (±0.065) 0.721 (±0.132) 0.624 (±0.041) U-m× (corr.) 0.524 (±0.023) 0.522 (±0.018) 0.621 (±0.154) 0.557 (±0.063) 0.629 (±0.038) 0.712 (±0.045) 0.437 (±0.133) 0.528 (±0.110) U-m× (gray) 0.547 (±0.062) 0.549 (±0.064) 0.742 (±0.145) 0.619 (±0.034) 0.626 (±0.038) 0.767 (±0.120) 0.437 (±0.175) 0.519 (±0.106)

### v2

@@ Results
Quality of Vulnerability Detection Datasets
[8] recently highlighted key challenges in existing vulnerability detection datasets, including label noise, data duplication, and data leakage. To mitigate these concerns, we incorporate the PrimeVul dataset in our experiments, as it was carefully curated to address such issues, along with several other commonly used datasets. We also include an additional unpublished dataset (NodeMedic) to demonstrate the generality of our approach. Beyond the noisiness of current vulnerability datasets, we acknowledge that function-level vulnerability detection has inherent limitations compared to repository-level detection [68]. However, we believe that function-level datasets and detectors provide a valuable first step toward addressing vulnerability detection in broader contexts.
VI Conclusion
Acknowledgment
This work was supported in part by the National Science Foundation via grants No. 2504353 and 2504354, and by the Future Enterprise Security Initiative at Carnegie Mellon CyLab (FutureEnterprise@CyLab).
Appendix A Full Results with Standard Deviation
TABLE III: Full results with standard deviation for all approaches across three models. Each cell reports the mean and standard deviation over five runs for Gemma and Qwen models, while GPT results are based on a single run and do not have standard deviation. For the approach names, the following abbreviations are used: ZS = zero-shot, R-FS = random few-shot, LFM = Learn-from-Mistakes, LFNN = Learn-from-Nearest-Neighbors, CM = Combined Method. A dash (‘–’) indicates that the metric is undefined in at least one of the runs due to division by zero. Dataset Approach Gemma-3-4b-it GPT-5-mini Qwen-2.5-Coder Acc Prec Recall F1 Acc Prec Recall F1 Acc Prec Recall F1 DiverseVul ZS 0.619 (±0.003) 0.770 (±0.009) 0.340 (±0.000) 0.472 (±0.002) 0.593 0.588 0.627 0.607 0.497 (±0.000) 0.000 (±0.000) 0.000 (±0.000) 0.000 (±0.000) R-FS 0.545 (±0.053) 0.571 (±0.081) 0.467 (±0.210) 0.480 (±0.132) 0.560 0.598 0.367 0.455 0.599 (±0.045) 0.741 (±0.027) 0.303 (±0.139) 0.411 (±0.150) LFM 0.525 (±0.003) 0.840 (±0.056) 0.061 (±0.005) 0.114 (±0.009) 0.617 0.647 0.513 0.573 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.512 (±0.004) 0.511 (±0.004) 0.537 (±0.009) 0.524 (±0.006) 0.590 0.624 0.453 0.525 0.659 (±0.009) 0.632 (±0.007) 0.759 (±0.014) 0.690 (±0.009) CM 1 0.548 (±0.020) 0.530 (±0.015) 0.868 (±0.035) 0.658 (±0.007) 0.587 0.610 0.480 0.537 0.638 (±0.008) 0.606 (±0.011) 0.792 (±0.039) 0.686 (±0.011) CM 2 0.515 (±0.002) 0.508 (±0.001) 0.973 (±0.000) 0.667 (±0.001) 0.560 0.573 0.473 0.518 0.597 (±0.006) 0.557 (±0.003) 0.949 (±0.008) 0.702 (±0.004) CM 3 0.499 (±0.002) 0.499 (±0.001) 0.985 (±0.007) 0.663 (±0.002) 0.570 0.585 0.480 0.527 0.531 (±0.021) 0.517 (±0.012) 0.988 (±0.015) 0.678 (±0.007) NodeMedic ZS 0.506 (±0.003) 0.765 (±0.005) 0.460 (±0.000) 0.574 (±0.001) 0.757 0.786 0.912 0.845 0.375 (±0.009) 0.870 (±0.049) 0.162 (±0.005) 0.273 (±0.008) R-FS 0.687 (±0.060) 0.726 (±0.003) 0.914 (±0.136) 0.804 (±0.060) 0.794 0.845 0.876 0.860 0.632 (±0.060) 0.748 (±0.020) 0.750 (±0.163) 0.739 (±0.071) LFM 0.725 (±0.000) 0.725 (±0.000) 1.000 (±0.000) 0.840 (±0.000) 0.767 0.789 0.927 0.852 0.720 (±0.000) 0.723 (±0.000) 0.993 (±0.000) 0.837 (±0.000) LFNN 0.758 (±0.004) 0.759 (±0.003) 0.975 (±0.004) 0.854 (±0.002) 0.788 0.849 0.861 0.855 0.713 (±0.011) 0.768 (±0.004) 0.866 (±0.014) 0.814 (±0.008) CM 1 0.701 (±0.005) 0.755 (±0.009) 0.870 (±0.018) 0.808 (±0.004) 0.788 0.839 0.876 0.857 0.751 (±0.008) 0.776 (±0.004) 0.923 (±0.020) 0.843 (±0.007) CM 2 0.565 (±0.008) 0.719 (±0.005) 0.657 (±0.008) 0.686 (±0.006) 0.794 0.840 0.883 0.861 0.735 (±0.006) 0.748 (±0.004) 0.958 (±0.003) 0.840 (±0.003) CM 3 0.722 (±0.007) 0.728 (±0.003) 0.984 (±0.014) 0.837 (±0.005) 0.794 0.846 0.883 0.864 0.720 (±0.000) 0.724 (±0.001) 0.991 (±0.003) 0.837 (±0.000) PrimeVul ZS 0.599 (±0.044) 0.723 (±0.106) 0.358 (±0.036) 0.472 (±0.003) 0.535 0.524 0.750 0.617 0.500 (±0.000) 0.002 (±0.004) 0.004 (±0.008) R-FS 0.522 (±0.041) 0.553 (±0.086) 0.373 (±0.148) 0.422 (±0.101) 0.560 0.556 0.600 0.577 0.503 (±0.010) 0.528 (±0.058) 0.232 (±0.105) 0.304 (±0.094) LFM 0.516 (±0.013) 0.726 (±0.186) 0.437 (±0.459) 0.336 (±0.270) 0.550 0.557 0.490 0.521 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.527 (±0.030) 0.520 (±0.017) 0.648 (±0.147) 0.571 (±0.064) 0.555 0.549 0.620 0.582 0.483 (±0.009) 0.390 (±0.056) 0.060 (±0.009) 0.104 (±0.015) CM 1 0.543 (±0.032) 0.527 (±0.022) 0.885 (±0.040) 0.659 (±0.010) 0.575 0.577 0.600 0.588 0.473 (±0.007) 0.462 (±0.010) 0.336 (±0.092) 0.383 (±0.057) CM 2 0.516 (±0.001) 0.508 (±0.001) 0.964 (±0.022) 0.666 (±0.005) 0.555 0.549 0.610 0.578 0.455 (±0.012) 0.445 (±0.015) 0.370 (±0.026) 0.404 (±0.021) CM 3 0.500 (±0.002) 0.500 (±0.001) 0.982 (±0.017) 0.663 (±0.004) 0.555 0.546 0.650 0.594 0.499 (±0.006) 0.499 (±0.003) 0.986 (±0.015) 0.663 (±0.005) SVENC ZS 0.478 (±0.000) 0.474 (±0.000) 0.400 (±0.000) 0.434 (±0.000) 0.533 0.523 0.756 0.618 0.500 (±0.000) 0.000 (±0.000) 0.000 (±0.000) R-FS 0.506 (±0.010) 0.503 (±0.005) 0.778 (±0.347) 0.564 (±0.174) 0.600 0.579 0.733 0.647 0.491 (±0.013) 0.422 (±0.373) 0.345 (±0.274) LFM 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) 0.544 0.543 0.556 0.549 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.471 (±0.009) 0.428 (±0.034) 0.187 (±0.044) 0.258 (±0.049) 0.567 0.554 0.689 0.614 0.504 (±0.009) 0.567 (±0.133) 0.040 (±0.009) 0.075 (±0.017) CM 1 0.520 (±0.008) 0.513 (±0.005) 0.813 (±0.027) 0.629 (±0.010) 0.578 0.566 0.667 0.612 0.520 (±0.019) 0.527 (±0.024) 0.387 (±0.074) 0.442 (±0.055) CM 2 0.536 (±0.004) 0.520 (±0.003) 0.947 (±0.011) 0.671 (±0.002) 0.589 0.574 0.689 0.626 0.493 (±0.005) 0.495 (±0.004) 0.636 (±0.011) 0.556 (±0.005) CM 3 0.500 (±0.000) 0.500 (±0.000) 0.938 (±0.029) 0.652 (±0.007) 0.611 0.596 0.689 0.639 0.496 (±0.005) 0.498 (±0.003) 0.982 (±0.009) 0.661 (±0.004) SVENP ZS 0.587 (±0.006) 0.597 (±0.007) 0.537 (±0.013) 0.565 (±0.009) 0.763 0.727 0.842 0.780 0.705 (±0.011) 0.965 (±0.028) 0.426 (±0.020) 0.591 (±0.019) R-FS 0.558 (±0.038) 0.561 (±0.044) 0.632 (±0.117) 0.584 (±0.035) 0.789 0.806 0.763 0.784 0.616 (±0.027) 0.707 (±0.071) 0.432 (±0.144) 0.514 (±0.102) LFM 0.500 (±0.000) 0.500 (±0.000) 0.905 (±0.013) 0.644 (±0.003) 0.776 0.744 0.842 0.790 0.500 (±0.000) 0.500 (±0.000) 1.000 (±0.000) 0.667 (±0.000) LFNN 0.634 (±0.005) 0.641 (±0.004) 0.611 (±0.011) 0.625 (±0.007) 0.895 0.941 0.842 0.889 0.792 (±0.005) 0.845 (±0.002) 0.716 (±0.011) 0.775 (±0.007) CM 1 0.553 (±0.012) 0.546 (±0.009) 0.616 (±0.043) 0.579 (±0.023) 0.882 0.892 0.868 0.880 0.787 (±0.023) 0.793 (±0.034) 0.779 (±0.021) 0.785 (±0.019) CM 2 0.632 (±0.000) 0.639 (±0.000) 0.605 (±0.000) 0.622 (±0.000) 0.895 0.917 0.868 0.892 0.697 (±0.019) 0.660 (±0.016) 0.816 (±0.029) 0.729 (±0.018) CM 3 0.521 (±0.013) 0.512 (±0.008) 0.942 (±0.045) 0.663 (±0.009) 0.803 0.795 0.816 0.805 0.600 (±0.018) 0.558 (±0.012) 0.963 (±0.027) 0.707 (±0.010)
An extended version of Table I with standard deviation is shown in Table III. A key observation is the general consistency of the methods—most exhibit low standard deviations. Table  III reinforces the reliability of the mean results discussed in the main text.